You have 775 mL of an 0.15 M acetic acid solution. What volume (V) of 1.80 M NaOH solution must you add in order to prepare an acetate buffer of pH = 4.65? (The pKa of acetic acid is 4.76.)
Answer #1

pH = 4.65

moles of acetic acid = 775 x 0.15 / 1000 = 0.116

moles of NaOH = 1.80 x V

CH3COOH +    NaOH   -------------------> CH3COONa + H2O

    0.11625            1.80V                                        0                 0

0.11625 - 1.80V      0                                         1.80 V         1.80 V

pH = pKa + log [salt / acid]

4.65 = 4.76 + log [1.80V / 0.11625 - 1.80 V]

[1.80V / 0.11625 - 1.80 V] = 0.776

V = 0.0282 L

volume of NaOH = 28.2 mL